# Elliptic asymptotic representation of the fifth Painlevé transcendents (corrected version)

### v8

@@ 2.2. Error term $\Delta(x)$
The elliptic expression above apparently contains the single integration constant $x_{0}$, and the other one is hidden in the error term $\Delta(x)\ll x^{-\delta}.$ To give a representation of $\Delta(x)$ we describe some necessary functions and constants.
For the same monodromy data $M^{0},$ $M^{1}$ as in Theorem 2.1, let $b(x)$ related to $y(x)$ be defined by
$a_{\phi}=A_{\phi}+\frac{B_{\phi}(t)}{t}=A_{\phi}+\frac{b(x)}{x}\quad\text{with $x=e^{i\phi}t$, \,\,\, i.e. $b(e^{i\phi}t)=e^{i\phi}B_{\phi}(t)$}$
(cf. (3.10), (7.6)). Set
@@
with $\vartheta^{\prime}(z,\tau)=(d/dz)\vartheta(z,\tau)$ is the $\vartheta$-function (cf. Section 5.2). Then $\psi_{0}(x)$ solves $2\psi^{\prime}_{0}=w(A_{\phi},\psi_{0})=\sqrt{(1-\psi_{0}^{2})(A_{\phi}-\psi_{0}^{2})},$ and $b_{0}(x)$ fulfills $b_{0}(e^{i\phi}t)-e^{i\phi}B_{\phi}(t)\ll t^{-\delta}$ in $S(\phi,t_{\infty},\kappa_{0},\delta_{0})$ (Proposition 5.6 and Corollary 6.1). Furthermore as in Section 7, $b^{\prime}_{0}(x)=2(\psi_{0}(x)^{2}-A_{\phi})+4\psi_{0}^{\prime}(x)$. Write
@@
in $S(\phi,t_{\infty},\kappa_{0},\delta_{0})$ [40, Theorem 2.1]. For (PI) or (PII) the relation between the error term $h(x)$ and $b(x)-b_{0}(x)$ in the $\tau$-function is referred to by Kitaev [28, p. 121].
@@ 3.1. Parametrisation of $y(x)$ by the monodromy data
[5, (2.8), (2.13)]. For the monodromy matrices $M^{0}=(m^{0}_{ij}),$ $M^{1}=(m^{1}_{ij})$, let $\mathcal{M}$ be the algebraic variety consisting of $(M^{0},M^{1})\in SL_{2}(\mathbb{C})^{2}$ such that
(3.2) $\mathrm{tr}\,M^{0}=2\cos\pi\theta_{0},\quad\mathrm{tr}\,M^{1}=2\cos\pi\theta_{1},\quad(M^{1}M^{0})_{11}=m^{1}_{11}m^{0}_{11}+m^{1}_{12}m^{0}_{21}=e^{-\pi i\theta_{\infty}},$
which is called the manifold of monodromy data. By (3.2) $\dim_{\mathbb{C}}\mathcal{M}=3.$
For a diagonal matrix $d_{0}^{\sigma_{3}}$ with $d_{0}\in\mathbb{C}\setminus\{0\}$, the gauge transformation $\Xi=d_{0}^{\sigma_{3}}\hat{\Xi}$ changes (1.1) into a system with $(y,\mathfrak{z},d_{0}^{-2}u)$ in place of $(y,\mathfrak{z},u)$, and the monodromy matrices for the matrix solution $\hat{\Xi}(\xi)$ of the same asymptotic form as of (2.1) become $d_{0}^{-\sigma_{3}}M^{0}d_{0}^{\sigma_{3}},$ $d_{0}^{-\sigma_{3}}M^{1}d_{0}^{\sigma_{3}}.$ By this fact combined with the surjectivity of the Riemann-Hilbert correspondence [5, §§3, 4, 5], [2, 3, 4, 7, 34], and the uniqueness [5, Propositions 2.1 and 2.2] (see also [12, Proposition 5.9 and Theorem 5.5], [13]) we have the following.
@@
Let $\mathcal{Y}(\mathrm{P}_{\mathrm{V}})$ be the family of solutions of $(\mathrm{P}_{\mathrm{V}})$, and let $\varphi:\,\mathcal{Y}(\mathrm{P}_{\mathrm{V}})\to\mathcal{M}$ be a map such that the image of each solution of $(\mathrm{P}_{\mathrm{V}})$ is the corresponding monodromy data. For $(M^{0},M^{1}),$ $(\tilde{M}^{0},\tilde{M}^{1})\in\mathcal{M},$ write $(M^{0},M^{1})\sim(\tilde{M}^{0},\tilde{M}^{1})$ if there exists $d_{0}\in\mathbb{C}\setminus\{0\}$ such that $(\tilde{M}^{0},\tilde{M}^{1})=d_{0}^{-\sigma_{3}}(M^{0},M^{1})d_{0}^{\sigma_{3}}$. Then we have the canonical bijection
$\varphi^{*}:\,\,\,\mathcal{Y}^{*}(\mathrm{P}_{\mathrm{V}})=\mathcal{Y}(\mathrm{P}_{\mathrm{V}})\setminus\mathcal{Y}_{0}(\mathrm{P}_{\mathrm{V}})\to\mathcal{M}^{*}=(\mathcal{M}\setminus\mathcal{M}_{0})/\sim,$
where
$\mathcal{M}_{0}=\{(M^{0},M^{1})\in\mathcal{M}\,|\,\,\,M^{0}=\pm I\,\,\,\text{or}\,\,\,M^{1}=\pm I\},$
$\mathcal{Y}_{0}(\mathrm{P}_{\mathrm{V}})=\{y\in\mathcal{Y}(\mathrm{P}_{\mathrm{V}})\,|\,\,\,\varphi(y)\in\mathcal{M}_{0}\}$
and $\dim_{\mathbb{C}}\mathcal{M}^{*}=2.$
Thus the solutions in $\mathcal{Y}^{*}(\mathrm{P}_{\mathrm{V}})$ are parametrised by $(M^{0},M^{1})\in\mathcal{M}^{*},$ essentially by two parameters.
Remark 3.1.
Instead of $(M_{0},M_{1})$ as above, i.e., a pair of the monodromy matrices for the solution (2.1), we may alternatively define $(M_{0},M_{1})$ as a conjugate class for the relation $(M_{0},M_{1})\sim T^{-1}(M_{0},M_{1})T$ with some $T\in SL_{2}(\mathbb{C})$ and $\mathcal{M}$ as the family of them. Then the bijection $\varphi^{*}$ becomes $\mathcal{Y}^{*}(\mathrm{P}_{\mathrm{V}})=\mathcal{Y}(\mathrm{P}_{\mathrm{V}})\setminus\mathcal{Y}_{0}(\mathrm{P}_{\mathrm{V}})\to\mathcal{M}^{*}=\mathcal{M}\setminus\mathcal{M}_{0}.$
@@ 3.2. Symmetric linear system
To consider $y(x)\in\mathcal{Y}^{*}(\mathrm{P}_{\mathrm{V}})$ along a ray $\arg x=\phi$ with $|\phi|<\pi/2,$ and to convert (1.1) to a symmetric form, set
@@
(ii) the isomonodromy deformation yields the same monodromy data $M^{0},$ $M^{1},$ $S_{1},$ $S_{2}$ as in Section 2, where $M^{0}$ and $M^{1}$ are given by the analytic continuation of $Y(t,\lambda)$ along $\hat{l}_{0}$ and $\hat{l}_{1}$, respectively, and $S_{k}$ are such that $Y_{k+1}(t,\lambda)=Y_{k}(t,\lambda)S_{k}$ with $Y_{k}(t,\lambda)$ ($Y_{2}(t,\lambda)=Y(t,\lambda)$) having the same asymptotic representation as (3.6) in the sector $|\arg\lambda-\pi/2-(k-2)\pi|<\pi$;
@@
$(y_{t}=dy/dt)$, and then $y(e^{i\phi}t)\in\mathcal{Y}^{*}(\mathrm{P}_{\mathrm{V}})$ is parametrised by $(M^{0},M^{1})\in\mathcal{M}^{*}.$
@@ 3.3. Characteristic roots, turning points and Stokes curves
In the remaining part of this paper, Sections 4 and 5 we are concerned with the direct monodromy problem for system (3.4), that is, calculation of the monodromy. Then $y,$ $\mathfrak{z}$ and $u$ may be treated as arbitrary complex parameters. It is also possible, in place of $\mathfrak{z}$, to choose $y^{*}=y_{t}$ related to $\mathfrak{z}$ via (3.8), and then let us suppose that $y,$ $y^{*}$ and $u$ are arbitrary complex parameters.
@@
as $t\to\infty$ outside the exceptional set $\bigcup_{\sigma\in Z(1)}\{|e^{i\phi}t-\sigma|<\delta_{1}\}$, $Z(1)=\{\sigma;y(\sigma)=1\}.$ Note that this is also valid in the cheese-like strip containing the ray $\arg x=\phi$:
$S^{\prime}(\phi,t^{\prime}_{\infty},\kappa_{0},\delta_{1})=\{e^{i\phi}t\,;\,\,\mathrm{Re\,}t>t^{\prime}_{\infty},\,|\mathrm{Im\,}t|<\kappa_{0}\}\setminus\bigcup_{\sigma\in Z(1)}\{|e^{i\phi}t-\sigma|\leq\delta_{1}\},$
$t^{\prime}_{\infty}$ and $\kappa_{0}$ $(>2\delta_{1})$ being given numbers. By (3.9), the turning points, that is, the zeros of $\mu$ are given by the following.
@@
On the positive real axis all the solutions of (PV) corresponding to the monodromy data such that $m^{0}_{11}m^{0}_{21}m^{1}_{11}m^{1}_{12}\not=0$ are given by [5, Theorems 3.1 and 4.1]. By using the expressions of these solutions it is easy to verify $a_{0}(t)$ $(=a_{\phi}(t)|_{\phi=0})\ll t^{-\varepsilon}$ as $t\to\infty$ (for solutions of [5, Theorem 4.1], as $t\to\infty$ along a suitable path avoiding poles). Then $\mathrm{Re\,}a_{\phi}(t)^{1/2}\ll|t^{-\varepsilon}|+|\phi|$ uniformly in $t$ for sufficiently small $|\phi|$, which implies that, as long as $m^{0}_{11}m^{0}_{21}m^{1}_{11}m^{1}_{12}\not=0$, every corresponding solution fulfills $0\leq\mathrm{Re\,}a_{\phi}(t)^{1/2}<1$ if $|\phi|$ is sufficiently small. On the other hand, for a general solution in [38, Theorem 2.18] along the imaginary axis, $a_{\pi/2}(t)=1+O(t^{-\varepsilon}).$
@@
To the monodromy data such that $m^{0}_{11}m^{0}_{21}m^{1}_{11}m^{1}_{12}=0$ correspond truncated solutions in sectors containing the positive real axis [2], [4]. Now, for these solutions, let us consider $a_{\phi}(t)$ with $(y,y^{*})=(y(t),y_{t}(t))$. Then we have $a_{\phi}(t)\ll t^{-1}$ for $\phi$ in some intervals containing $\phi=0$. In the case $m^{1}_{11}=0$ the solution $y(x)\sim-1+cx^{-1/2}e^{ix/2}$ in $0\leq\arg x\leq\pi$ [4, Proposition 5], [5, Corollary 5.2] satisfies $a_{\phi}(t)\ll t^{-1}$ for $0\leq\phi\leq\pi.$ If $m^{0}_{11}=m^{1}_{11}=0,$ then $y(x)\sim-1+4(\theta_{0}+\theta_{1}-1)x^{-1}$ in $|\arg x|<\pi$ [4, Proposition 2], [5, Corollary 5.3] satisfies $a_{\phi}(t)\ll t^{-2}$ for $|\phi|<\pi.$ For $m^{1}_{12}=0$ or $m^{0}_{21}=0$, truncated solutions such that $a_{\phi}(t)\ll t^{-1}$ for $|\phi|<\pi/2$ are given by [2].
@@
as $t\to\infty.$ Here $g_{2}(t,\lambda)$ has branch points at $\lambda^{0}_{1,2},$ $\lambda_{1,2}$, $\pm e^{i\phi}$ and $\pm e^{i\phi}a_{\phi}^{1/2}$, but it fulfills $g_{2}(t,\lambda)\ll 1$ if $|\lambda^{2}-e^{2i\phi}a_{\phi}|^{-1}+|\lambda^{2}-e^{2i\phi}|^{-1}\ll 1.$ The algebraic function $\mu(t,\lambda)$ is given on the Riemann surface consisting of two copies of $\lambda$-plane $\mathbb{P}_{+}$ and $\mathbb{P}_{-}$ glued along the cuts $\lambda_{1},\lambda^{0}_{1}], [$$\lambda^{0}_{2},\lambda_{2}]$ (cf. Figure 3.2, (a)). Note that, in (3.11), each square root is fixed in such a way that
$a^{-1/2}_{\phi}\sqrt{\frac{e^{2i\phi}a_{\phi}-\lambda^{2}}{e^{2i\phi}-\lambda^{2}}},\quad e^{-2i\phi}a^{-1/2}_{\phi}\sqrt{(e^{2i\phi}a_{\phi}-\lambda^{2})(e^{2i\phi}-\lambda^{2})}\to\pm 1$
as $\lambda\to 0$ on $\mathbb{P}_{\pm},$ $a_{\phi}^{1/2}$ being as in Proposition 3.2.

### v9

@@ 2.2. Error term $\Delta(x)$
The elliptic expression above apparently contains the single integration constant $x_{0}$, and the other one is hidden in the error term $\Delta(x)\ll x^{-\delta}.$ Let us define necessary functions and constants to express $\Delta(x)$.
For $y(x)=y(x,M^{0},M^{1})$ of Theorem 2.1, let $b(x)$ be such that
$a_{\phi}=A_{\phi}+\frac{B_{\phi}(t)}{t}=A_{\phi}+\frac{b(x)}{x}\quad\text{with $x=e^{i\phi}t$, \,\,\, i.e. $b(e^{i\phi}t)=e^{i\phi}B_{\phi}(t)$},$
which is related to $y(x)$ via (3.10), (7.5) with $y^{*}=y^{\prime}(t)$. Set
@@
with $\vartheta^{\prime}(z,\tau)=(d/dz)\vartheta(z,\tau)$ is the $\vartheta$-function (cf. Section 5.2). Then $\psi_{0}(x)$ solves $2\psi^{\prime}_{0}=w(A_{\phi},\psi_{0})=\sqrt{(1-\psi_{0}^{2})(A_{\phi}-\psi_{0}^{2})},$ and $b_{0}(x)$ fulfills $b_{0}(x)-b(x)=b_{0}(e^{i\phi}t)-e^{i\phi}B_{\phi}(t)\ll t^{-\delta}$ in $S(\phi,t_{\infty},\kappa_{0},\delta_{0})$ (Proposition 5.6 and Corollary 6.1). Furthermore as in Section 7, $b^{\prime}_{0}(x)=2(\psi_{0}(x)^{2}-A_{\phi})+4\psi_{0}^{\prime}(x)$. Write
@@
in $S(\phi,t_{\infty},\kappa_{0},\delta_{0})$ [40, Theorem 2.1].
@@ 3.1. Parametrisation of $y(x)$ by the monodromy data
[5, (2.8), (2.13)]. For the monodromy matrices $M^{0}=(m^{0}_{ij}),$ $M^{1}=(m^{1}_{ij})$, let $\mathcal{M}^{*}_{(\theta_{0},\theta_{1},\theta_{\infty})}$ be the algebraic variety consisting of $(M^{0},M^{1})\in SL_{2}(\mathbb{C})^{2}$ such that
(3.2) $\begin{split}&\mathrm{tr}\,M^{0}=2\cos\pi\theta_{0},\quad\mathrm{tr}\,M^{1}=2\cos\pi\theta_{1},\\ &(M^{1}M^{0})_{11}=m^{1}_{11}m^{0}_{11}+m^{1}_{12}m^{0}_{21}=e^{-\pi i\theta_{\infty}},\end{split}$
which is called the manifold of monodromy data in [5]. Then $\dim_{\mathbb{C}}\mathcal{M}^{*}_{(\theta_{0},\theta_{1},\theta_{\infty})}=3.$ Our monodromy manifold $\mathcal{M}_{(\theta_{0},\theta_{1},\theta_{\infty})}$ defined in Section 2 is written as $\mathcal{M}_{(\theta_{0},\theta_{1},\theta_{\infty})}=\mathcal{M}^{*}_{(\theta_{0},\theta_{1},\theta_{\infty})}/\sim$. For $d_{0}\in\mathbb{C}\setminus\{0\}$, the gauge transformation $\Xi=d_{0}^{\sigma_{3}}\hat{\Xi}$ changes (1.1) with $(y,\mathfrak{z},u)$ to an isomonodromy system with $(y,\mathfrak{z},d_{0}^{-2}u)$, and the monodromy matrices for the canonical solution $\hat{\Xi}(\xi)$ become $d_{0}^{-\sigma_{3}}M^{0}d_{0}^{\sigma_{3}},$ $d_{0}^{-\sigma_{3}}M^{1}d_{0}^{\sigma_{3}}.$ By this fact combined with the surjectivity of the Riemann-Hilbert correspondence [5, §§3, 4, 5], [2, 3, 4, 7, 34], and the uniqueness [5, Propositions 2.1 and 2.2] (see also Proposition 3.2, [12, Proposition 5.9 and Theorem 5.5], [13]) we have the following.
@@
Let $\mathcal{Y}(\mathrm{P}_{\mathrm{V}})$ be the family of solutions of $(\mathrm{P}_{\mathrm{V}})$, and let $\varphi:\,\mathcal{Y}(\mathrm{P}_{\mathrm{V}})\to\mathcal{M}_{(\theta_{0},\theta_{1},\theta_{\infty})}$ be such that $y\mapsto(M^{0},M^{1})$ if $(y,\mathfrak{z},u)$ governs the isomonodromy deformation of (1.1) with monodromy data $(M^{0},M^{1})$. Then we have the canonical bijection
$\varphi:\,\,\,\mathcal{Y}(\mathrm{P}_{\mathrm{V}})\setminus\mathcal{Y}_{0}(\mathrm{P}_{\mathrm{V}})\to\mathcal{M}_{(\theta_{0},\theta_{1},\theta_{\infty})}\setminus\mathcal{M}_{0},$
where $\mathcal{M}_{0}=\{(M^{0},M^{1})\in\mathcal{M}_{(\theta_{0},\theta_{1},\theta_{\infty})}\,|\,\,M^{0}=\pm I\,\,\text{or}\,\,M^{1}=\pm I\}$, $\mathcal{Y}_{0}(\mathrm{P}_{\mathrm{V}})=\varphi^{-1}(\mathcal{M}_{0}).$
Thus the solutions in $\mathcal{Y}(\mathrm{P}_{\mathrm{V}})\setminus\mathcal{Y}_{0}(\mathrm{P}_{\mathrm{V}})$ are parametrised by $(M^{0},M^{1})\in\mathcal{M}_{(\theta_{0},\theta_{1},\theta_{\infty})}\setminus\mathcal{M}_{0}$, essentially two parameters.
@@ 3.2. Symmetric linear system
To consider $y(x)\in\mathcal{Y}(\mathrm{P}_{\mathrm{V}})$ along a ray $\arg x=\phi$ with $|\phi|<\pi/2,$ and to convert (1.1) to a symmetric form, set
@@
(ii) the isomonodromy deformation yields the same monodromy data $M^{0},$ $M^{1},$ $S_{1},$ $S_{2}$ as in Section 2, where $M^{0}$ and $M^{1}$ are defined by $Y^{(\hat{l}_{\nu})}(t,\lambda)=Y(t,\lambda)M^{\nu}$ for $\nu=0,1$ with $Y^{(\hat{l}_{\nu})}(t,\lambda)$ denoting the analytic continuation of $Y(t,\lambda)$ along the loop $\hat{l}_{\nu}$, and $S_{k}$ are such that $Y_{k+1}(t,\lambda)=Y_{k}(t,\lambda)S_{k}$ with $Y_{k}(t,\lambda)$ ($Y_{2}(t,\lambda)=Y(t,\lambda)$) having the same asymptotic representation as (3.6) in the sector $|\arg\lambda-\pi/2-(k-2)\pi|<\pi$;
@@
$(y_{t}=dy/dt)$, and then $y(e^{i\phi}t)$ is parametrised by $(M^{0},M^{1})\in\mathcal{M}_{(\theta_{0},\theta_{1},\theta_{\infty})}.$
@@
The uniqueness of the Riemann-Hilbert correspondence for system (3.4) will be used in the justification procedure in Section 6.
Proposition 3.2.
If $M^{0}\not=I$ and $M^{1}\not=I$, a canonical solution $Y_{2}(t,\lambda)$ of (3.4) corresponding to $(M^{0},M^{1})$ is uniquely determined.
Proof..
Let $Y_{2}(\lambda)$ and $\tilde{Y}_{2}(\lambda)$ be canonical solutions defining $(M^{0},M^{1}).$ Suppose that $Y_{2}(\lambda)$ and $\tilde{Y}_{2}(\lambda)$ admit the same asymptotic representation (3.6) as $\lambda\to\infty$ in the sector $|\arg\lambda-\pi/2|<\pi,$ and set, around $\lambda=-e^{i\phi}$, i.e., $\lambda_{-}:=\lambda+e^{i\phi}=0,$
$Y_{2}(\lambda)=G_{0}(I+O(\lambda_{-}))\lambda_{-}^{(\theta_{0}/2)\sigma_{3}+L}C_{0},\quad\tilde{Y}_{2}(\lambda)=\tilde{G}_{0}(I+O(\lambda_{-}))\lambda_{-}^{(\theta_{0}/2)\sigma_{3}+\tilde{L}}\tilde{C}_{0},$
where $G_{0},$ $\tilde{G}_{0},$ $C_{0},$ $\tilde{C}_{0}\in SL_{2}(\mathbb{C})$ and $L=(l_{ij})$ and $\tilde{L}=(\tilde{l}_{ij})$ are upper-triangular (respectively, lower-triangular) matrices with vanishing diagonal entries. Then we have $M^{0}=C_{0}^{-1}e^{\pi i\theta_{0}\sigma_{3}}e^{2\pi iL}C_{0}=\tilde{C}_{0}^{-1}e^{\pi i\theta_{0}\sigma_{3}}e^{2\pi i\tilde{L}}\tilde{C}_{0},$ and hence $M^{0}\not=\pm I$ implies either of the cases: (i) $\theta_{0}\not\in\mathbb{Z}$ and $L=\tilde{L}=0;$ (ii) $\theta_{0}\in\mathbb{Z}_{\geq 0}$ and $l_{12}\tilde{l}_{12}\not=0$ (respectively, $\theta_{0}\in\mathbb{Z}_{\leq 0}$ and $l_{21}\tilde{l}_{21}\not=0$). The relation $(C_{0}\tilde{C}_{0}^{-1})^{-1}e^{\pi i\theta_{0}\sigma_{3}}e^{2\pi iL}C_{0}\tilde{C}_{0}^{-1}=e^{\pi i\theta_{0}\sigma_{3}}e^{2\pi i\tilde{L}}$ leads to $C_{0}\tilde{C}_{0}^{-1}=\mathrm{diag}[\alpha,\alpha^{-1}]$ or $\mathrm{diag}[\alpha,\alpha^{-1}]+L_{*}$ with some $\alpha\not=0$ and some upper-triangular (respectively, lower-triangular) matrix $L_{*}$ with vanishing diagonal entries. By this fact
$Y_{2}(\lambda)\tilde{Y}_{2}(\lambda)^{-1}=G_{0}(I+O(\lambda_{-}))\lambda_{-}^{(\theta_{0}/2)\sigma_{3}}e^{2\pi iL}C_{0}\tilde{C}_{0}^{-1}e^{-2\pi i\tilde{L}}\lambda_{-}^{-(\theta_{0}/2)\sigma_{3}}(I+(\lambda_{-}))\tilde{G}_{0}^{-1}$
is holomorphic around $\lambda=-e^{i\phi}$. Similarly $Y_{2}(\lambda)\tilde{Y}_{2}(\lambda)^{-1}$ is also holomorphic around $\lambda=e^{i\phi}$ if $M^{1}\not=\pm I.$ Observing that $Y_{2}(\lambda)\tilde{Y}_{2}(\lambda)^{-1}=I+O(\lambda^{-1})$ as $\lambda\to\infty$ through $|\arg\lambda-\pi/2|<\pi$, we conclude $Y_{2}(\lambda)=\tilde{Y}_{2}(\lambda)$ by the Phragmén-Lindelöf reasoning. ∎
@@ 3.3. Characteristic roots, turning points and Stokes curves
In the remaining part of this section, Sections 4 and 5 we are concerned with the direct monodromy problem for system (3.4), that is, calculation of the monodromy. Then $y,$ $\mathfrak{z}$ and $u$ may be treated as arbitrary complex parameters. It is also possible, in place of $\mathfrak{z}$, to choose $y^{*}=y_{t}$ related to $\mathfrak{z}$ via (3.8), and then let us suppose that $y,$ $y^{*}$ and $u$ are arbitrary complex parameters.
@@
By (3.9), as long as $y\not=0,1,\infty$ and $y^{*}\not$$\infty$, the turning points, that is, the zeros of $\mu$ are given by the following.
@@
For a solution $y(t)$ of (PV), let us consider $a_{\phi}(t)$ with $(y,y^{*})=(y(t),y_{t}(t))$. On the positive real axis all the solutions corresponding to the monodromy data such that $m^{0}_{11}m^{0}_{21}m^{1}_{11}m^{1}_{12}\not=0$ are given by [5, Theorems 3.1 and 4.1]. By using the expressions of these solutions it is easy to verify $a_{0}(t)$ $(=a_{\phi}(t)|_{\phi=0})\ll t^{-\varepsilon}$ as $t\to\infty$ (for solutions of [5, Theorem 4.1], as $t\to\infty$ along a suitable path avoiding poles). Then $\mathrm{Re\,}a_{\phi}(t)^{1/2}\ll|t^{-\varepsilon}|+o(1)$ uniformly in $t$ for sufficiently small $|\phi|$, which implies that, as long as $m^{0}_{11}m^{0}_{21}m^{1}_{11}m^{1}_{12}\not=0$, every corresponding solution fulfills $0\leq\mathrm{Re\,}a_{\phi}(t)^{1/2}<1$ if $|\phi|$ is sufficiently small. On the other hand, for a general solution in [38, Theorem 2.18] along the imaginary axis, $a_{\pi/2}(t)=1+O(t^{-\varepsilon}).$
@@
To the monodromy data such that $m^{0}_{11}m^{0}_{21}m^{1}_{11}m^{1}_{12}=0$ correspond truncated solutions in sectors containing the positive real axis [2], [4]. Let us consider $a_{\phi}(t)$ with $(y,y^{*})=(y(t),y_{t}(t))$. Then we have $a_{\phi}(t)\ll t^{-1}$ for $\phi$ in some intervals containing $\phi=0$. In the case $m^{1}_{11}=0$ the solution $y(x)\sim-1+cx^{-1/2}e^{ix/2}$ in $0\leq\arg x\leq\pi$ [4, Proposition 5], [5, Corollary 5.2] satisfies $a_{\phi}(t)\ll t^{-1}$ for $0\leq\phi\leq\pi.$ If $m^{0}_{11}=m^{1}_{11}=0,$ then $y(x)\sim-1+4(\theta_{0}+\theta_{1}-1)x^{-1}$ in $|\arg x|<\pi$ [4, Proposition 2], [5, Corollary 5.3] satisfies $a_{\phi}(t)\ll t^{-2}$ for $|\phi|<\pi.$ For $m^{1}_{12}=0$ or $m^{0}_{21}=0$, truncated solutions such that $a_{\phi}(t)\ll t^{-1}$ for $|\phi|<\pi/2$ are given by [2].
@@
as $t\to\infty.$ Here $g_{2}(t,\lambda)$ has branch points at $\lambda^{0}_{1,2},$ $\lambda_{1,2}$, $\pm e^{i\phi}$ and $\pm e^{i\phi}a_{\phi}^{1/2}$, but it fulfills $g_{2}(t,\lambda)\ll 1$ if $|\lambda^{2}-e^{2i\phi}a_{\phi}|^{-1}+|\lambda^{2}-e^{2i\phi}|^{-1}\ll 1.$ The algebraic function $\mu(t,\lambda)$ is given on the Riemann surface consisting of two copies of $\lambda$-plane $\mathbb{P}_{+}$ and $\mathbb{P}_{-}$ glued along the cuts $\lambda_{1},\lambda^{0}_{1}], [$$\lambda^{0}_{2},\lambda_{2}]$ (cf. Figure 3.2, (a)). In (3.11), each square root is fixed in such a way that
$\sqrt{\frac{e^{2i\phi}a_{\phi}-\lambda^{2}}{e^{2i\phi}-\lambda^{2}}}\to 1,\quad\lambda^{-2}\sqrt{(e^{2i\phi}a_{\phi}-\lambda^{2})(e^{2i\phi}-\lambda^{2})}\to-1$
as $\lambda\to\infty$ on $\mathbb{P}_{+}.$ Then, for $a_{\phi}^{1/2}$ as in Proposition 3.3, $a^{-1/2}_{\phi}\sqrt{(e^{2i\phi}a_{\phi}-\lambda^{2})/(e^{2i\phi}-\lambda^{2})}$, $e^{-2i\phi}a^{-1/2}_{\phi}\sqrt{(e^{2i\phi}a_{\phi}-\lambda^{2})(e^{2i\phi}-\lambda^{2})}\to 1$ as $\lambda\to 0$ on $\mathbb{P}_{+}$.